OpenStudy (anonymous):

Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita's score was 40 less than twice Joelle's. What were their scores?

OpenStudy (anonymous):

Set up the Equation: Anita Score = a, Joelle's score = j Together they score 425: a + j = 425 Anita's score was 40 less than twice Joelle a= 2j -40 Substitute 2j + j -40 =425 Solve for J 3j = 465 j=155 Plug in J to solve for A a=2j -40 a=2(155) -40 a= 310-40 a=270 Anita Scored 270 and Joelle score 155 Added together that does score 425